Wheeler’s Bike Company manufactures custom racing bicycles. The company uses a job order cost system to determine the cost of each bike. Estimated costs and expenses for the coming year follow: Bike parts $ 351,800 Factory machinery depreciation 54,500 Factory supervisor salaries 139,500 Factory direct labor 238,266 Factory supplies 38,900 Factory property tax 28,750 Advertising cost 22,500 Administrative salaries 47,000 Administrative-related depreciation 18,700 Total expected costs $ 939,916 Required:

1. Calculate the predetermined overhead rate per direct labor hour if the average direct labor rate is $12.81 per hour. (Round your answer to 2 decimal places.)

2. Determine the amount of applied overhead if 18,900 actual hours are worked in the upcoming year. (Round your intermediate calculations to 2 decimals. Round your final answer to the nearest whole dollar.)
Answer #1

1)

First we should determine expected number of DL hours.

Estimated DL Cost/DL rate = Estimated DL hours

238,266/12.81 = 18,600

Predetermined Overhead Rate = Estimated Mfg. Overhead / Estimated DL hours

Estimated Total Manufacturing Overhead:

Factory machinery depreciation 54,500
Factory supervisor salaries 139,500
Factory supplies 38,900
Factory property tax 28,750
Total Estimated MOH 261,650

Predetermined Overhead Rate = 261,650/18600

= 14.067(rounded)

2) Applied Overhead = Overhead Rate*Actual DL hours

14.067*18,900 = 265,866